OpenStudy (errrwut):

(Link to document inside) I do not understand the intermediate steps that were taken to solve for a constant of an integral.

OpenStudy (errrwut):

http://people.whitman.edu/~hundledr/courses/M367/HWSect01_4.pdf Problem 1.4.2 (B)

OpenStudy (errrwut):

Where they solve the ODE: \[u _{xx} = -\frac{ Q }{ K_{0} } = -x \rightarrow u_{x} = -\frac{1}{2}x ^{2}+\frac{1}{6}L^{2}\] I do not understand where they got: \[+\frac{1}{6}L^{2}\] Because if you solve the integral: \[\int\limits_{0}^{L}u_{xx} dx =\int\limits_{0}^{L}(-x) dx\] Don't you just get \[-\frac{x^{2}}{2}+C_{1}\] Right?

OpenStudy (errrwut):

I mean, if you solve for \[C_{1}\] You just get: \[C_{1} = \frac{L^{2}}{2}\] and \[u_{x} = -\frac{1}{2}x ^{2}+\frac{1}{2}L^{2}\] ?

zepdrix (zepdrix):

Hmm I don't remember PDE's very well :\ How are you solving for the constant? They gave us information about the boundary conditions for u, but not for u', so I'm a little confused. \(\large\rm u_x(L)=?\) \(\large\rm u_x(0)=?\)

OpenStudy (errrwut):

I am also confused. I was just using x = 0, x = L as the upper and lower bounds since it said to integrate the entire rod So, \[\int\limits_{0}^{L}u_{xx} dx =\int\limits_{0}^{L}(-x) dx \] Then \[u_{x} = -\frac{x^{2}}{2}+C_{1}|_{0}^{L}\] \[=-\frac{(L)^{2}}{2} + C_{1} - (-\frac{0^{2}}{2})\] \[C_{1}=\frac{L^{2}}{2}\] I think this might be the better way \[u_{x} =-\frac{x^{2}}{2}+C_{1}\] for \[ x = 0\] \[u_{0} =-\frac{(0)^{2}}{2}+C_{1}\] Then \[C_{1} = 0\] \[u_{x} =-\frac{x^{2}}{2}\] for \[ x = L\] \[u_{x} =-\frac{(L)^{2}}{2}+C_{1}\] Then \[C_{1} = \frac{L^{2}}{2}\] \[u_{x} =-\frac{x^{2}}{2} + \frac{L^{2}}{2}\]

OpenStudy (errrwut):

UPDATE: Solved at x = 0 and x = L \[\phi(x,t) = -K_{0}u_{x}(x,t)\] We have: ODE: \[u_{x}(x,t) = K_{0} x\] BC: \[u(0) = 0 \] and \[u(L) = 0\] Solving the ODE: \[\phi(x,t) = -K_{0}u_{x}(x,t)\] \[\int\limits_{}^{}u_{xx}dx = \int\limits_{}^{}K_{0}xdx\] \[u_{x} = \frac{K_{0}x^{2}}{2} + C_{1}\] \[\int\limits_{}^{}u_{x}dx = \int\limits_{}^{}\frac{K_{0}x^{2}}{2} + C_{1}dx\] \[ = \int\limits_{}^{}\frac{K_{0}x^{2}}{2}dx + \int\limits_{}^{} C_{1}dx\] \[= \frac{K_{0}}{2}\int\limits_{}^{}x^{2}dx + C_{1}x\] \[= \frac{K_{0}}{2}(\frac{x^3}{3}) + C_{1}x\] \[= \frac{K_{0}x^3}{6} + C_{1}x\] \[u(x,t) = \frac{K_{0}x^3}{6} + C_{1}x + C_{2}\] Apply boundary conditions (BC) to \[u(x,t)\] and solve for C1 and C2, and then take the derivative of u(x,t) so that you can use it in: \[\phi(x,t) = -K_{0}u_{x}(x,t)\] Where x = 0, and x = L